Reports of 56424
Detector Operation (Operations Report)
menzione - 16:00 Saturday 23 September 2023 (61758) Print this report
Operator Report - Morning shift

Upon arrival I found ITF locked at LN3 since 4h. Unfortunately it unlocked and I spent all the shift to try to acheive a stable lock.
The microseism due to the sea activity (fig1) prevented the lock for the whole shift. 
ITF locked in LN2 at 13:57 UTC. For 1 minute.
Etalon tuning not performed.

Sub-system reports

SBE
After the unlock of 06:00 UTC, SIB2 vertical position drifted. I recover it via stepping motors.

Oncall events

ISC
(23-09-2023 10:00 - 23-09-2023 12:00) From remote
Status: Ended
Description: ITF lock instability

Images attached to this report
AdV-COM (AdV commissioning (1st part) )
bersanetti - 13:04 Saturday 23 September 2023 (61759) Print this report
Lock (acquisition) stability

I gave a look at the unlocks from this morning, both during lock acquisition and in the "steady" states; besides a couple of them, they fall in three categories:

  • unlocks during the CARM handoff to the IMC: long standing story, they have been quite reduced in frequency after the last change in CARM offset at the time of the handoff, but it looks like that bad weather (as in wind and sea activity, which today are both annoying) has an impact on those;
  • unlocks at apparently steady state in CARM_NULL_1F: "apparently" as they happen just a few seconds after getting there; I could find one definite instance, and two other suspect ones, of the fact that they are induced by the engagement of the CARM2RFC loop, which either saturates (Figure 1) and/or pushes a lot with "fast" transients +-5/6 V (Figures 2 and 3); for the time being I tried using a different CARM filter (1 success out of 2 at the time of writing), I will observe during the day and maybe change strategy (lower gain to act more as a drift control to avoid fast/big corrections);
  • unlocks in LOW_NOISE_2: a couple of them happened (Figures 4 to 7, two unlocks) with a quickly diverging 55 Hz oscillation, most probably induced by the diagonalization process which abruptly changes sign to a couple of contributions; I disabled the re-engagement of PyDiag in LOW_NOISE_2 and will observe behaviour/performance.
Images attached to this report
AdV-DET (Commissioning)
gouaty, hui, nardecchia, pinto, magazzu, mwas, mantovani, tringali - 11:04 Saturday 23 September 2023 (61753) Print this report
OMC mode matching tuning versus CMRF

(We tried to post this entry during last night, it appeared on my local session of the logbook, but in fact the post did not work).

The goal of the afternoon shift was to find the longitudinal position of SDB1 (Z) that minimizes the CMRF.

 

At the beginning of the shift, Maddalena commented the engagement of the BS_TY_SET loop in ITF_LOCK.py.

Then there were several unsuccessful lock trials, two unlocks happened right after reaching CARM_NULL_1F. It was then decided to cancel out the last step performed on the NI DAS around 15h20 utc.

While the ITF was in LOCKED_ARMS_BEAT_DRMI_3F, we saw that the SDB1 bench was moving (the drift control was not working). The galvo loops of B1p QD were also open so we closed them manually at 15h18m42-UTC. After this action, the drift control of SDB1 was back in operation.

After several unsuccessful attempts to lock the ITF finally reached a stable lock in CARM_NULL_1F at 16h37 utc.

ITF locked in DC readout at 16h56m30 utc but we unlocked shortly after at 16h57m06 utc during the transition to Low Noise 1.

DC readout reached again at 17h32m10 utc, but ITF unlocked again at 17h34m14 utc during the transition to Low Noise 1. The ISC team noticed the presence of an oscillation on BS_TX at 275 mHz. For safety they are disabling the full bandwidth control of BS for the next relock.

Noticing that the B1p 56 MHz side bands were very unblanced, an offset was added on MICH to improve the balancing of B1p side bands.

It was then possible to reach DC readout at 18h50m05 utc. and Low Noise 2 at 18h52 utc but ITF unlocked at 18h55m18 utc.

DC readout at 19h57 utc and Low Noise 2 at 19h59 utc. Low Noise 3 reached at 20h26 utc.

The BS_TY_SSFS_LF_Q_ENBL was disabled when arriving in Low Noise 3 (because it had been renabled before for a few min), and the offsets were cancelled to bring BS_TY_INPUT back at 0.

 

We then moved the SDB1 bench along Z (starting from initial position: SDB1_LC_Z = 3890 um, Sa_OB_F0_X = 2599 um).

We performed steps of 100 um on F0_X, and we adjusted the bench TX offset in parallel by doing steps of -2 urad per 150 s, for each step of 100 um of bench displacement.

We stop at 22h43 utc at SDB1_LC_Z = 960 um, Sa_OB_F0_X = -900 um which seems to correspond to a zero on the B1s signal demodulated at 491 Hz. We take quiet data until 23h06 utc, unfortunately the DCP was only getting worse.

Fig.1 shows the scan of the SDB1 bench longitudinal position and the period of data when we stabilized it at SDB1_LC_Z = 960 um. Data will be analyzed offline.

Fig.2 shows the scan of SDB1 in the opposite direction towards the standard position. The ITF unlocked at 23h49 utc.

After the unlock we restored the bench at its standard position and re-enabled the BS_TY_SSFS_LF_Q loop in ITF_LOCK.

Images attached to this report
Detector Operation (Operations Report)
menzione - 7:53 Saturday 23 September 2023 (61757) Print this report
Operator Report - Initial status

ITF State: LN3
ITF Mode: LOCKED
Quick Summary: ISYS: in standard state, SBE loops closed, Susp loops closed, DSP of BS still offline.
Activities ongoing since this morning: No activities ongoing

Images attached to this report
Detector Operation (Operations Report)
amagazzu - 2:22 Saturday 23 September 2023 (61756) Print this report
Operator Report - Afternoon shift

The first part of the shift was dedicated to the recovery of the of the ITF up to LOW_NOISE_3 (see report #61755). We reached LOW_NOISE_2 at 14:02 UTC, the lock lasted until 14:21 UTC. Unfortunately after this event we were able to reach LOW_NOISE_3, required for the DET activity, only at 20:25 UTC. 
DET activity concluded at 1:50 UTC due to an unlock, while relocking Gouaty restored the old SDB1 position. ITF left relocking with AUTORELOCK_FAILSAFE engaged. 

Sub-system reports

DAQ
During the morning two blackout events, happened at around 10:00 UTC and 12:00 UTC, triggered some missing data from various ACS INJ and DET signals. After restarting the processes eurothermServer and BacnetServer and reloading FbsMoni at around 14:10 UTC the signals were achieved again. Also, thanks to the support of Cavalieri and Kraja, the module ENV-WAB was remotely restarted and the signal related to the WAB Building were restored from 15:15 UTC.

ISC
At 0:03 UTC, as agreed with the Commissioning crew, I changed the Etalon WI Setpoint from 20.25 to 20.35 with a ramp of 4 hours (command 'WI:WI Set point [WI_RH_SET=20.35,rampTime=14400]').

Images attached to this report
AdV-TCS (CO2 laser projector)
nardecchia, pinto, gouaty, magazzu, tringali - 22:51 Friday 22 September 2023 (61755) Print this report
Comment to DAS tuning (61754)

Today afternoon, it was not possible to reach LN2 most probably due to the unbalance of the B1p 56 MHz upper and lower sidebands (see fig.1).

In the lock started at 18.11 UTC (when CARM NULL 1F has been achieved), we added -11 as MICH offset in order to overimpose the sidebands  to proceed up to LN2, the trick worked (see fig.2) but the ITF unlocked at LN2 due to huge oscillation at ~55 Hz .  

Thus, to save the weekend, we decided to reverse the last step of DAS WP.

So, at 19.24 UTC, we came back to STEP_1S (61729) .

 

STEP_1S
19.24 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.090

1.857

on the pickoff

0.418

 0.0146

0.302

N

on the ITF

0.091

0.266 

2.972

on the pickoff

0.561

0.044

0.484

After that, the ITF achieved LN3 (at 20.25 UTC) without adding extra offsets (fig.3).

As requested by Romain, the BS TY SET was put to 0.

We left the ITF to Romain.

Images attached to this comment
AdV-TCS (CO2 laser projector)
mantovani, nardecchia, pinto, ruggi, spinicelli - 17:48 Friday 22 September 2023 (61754) Print this report
DAS tuning

After solving the problem on the BS DSP, DAS tuning started.

9.10 UTC: LN3

9.35 UTC: WI DAS INCREASE (+15 %)

10.01UTC: unlock

10.53 UTC: the sidebands was low in the CITF preventing the lock acquisition→ increase WI CH by 15 %

The powers applied in this step are summarized in the table below.

 

STEP6_S

12.39 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.078

0.090*(1+15/100)=0.119

1.615*(1+15/100)=2.45

on the pickoff

0.48

0.0169*(1+15/100)= 0.0194

0.347*(1+15/100)=0.399

N

on the ITF

0.091

0.266 

2.972

on the pickoff

0.561

0.044

0.484

After that, it was possible to reach CARM NULL 1F but the misbalance of the upper and lower 56 MHz sidebands on B1p prevented the lock in low noise, thus I reverted the increase of the WI DAS power and, following Paolo’s indication, I slightly decreased the NI DAS by 10 % (at 13.28 UTC).

 

STEP7_S
13.28 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.078

0.104

2.13

on the pickoff

0.48

0.0169

0.347

N

on the ITF

0.091

0.186*(1-10/100)+0.08=

0.247 

1.172*(1-10/100)+1.8 =2.85

on the pickoff

0.561

0.036*(1-10/100)+0.013=

0.0402

0.190*(1-10/100)+0.293=

0.464

 

The ITF unlocked during different steps of the lock acquisition, thus, the DAS powers have been re-settled to STEP 5S (at 15.22 UTC).

 

STEP5_S 

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.104

2.13

on the pickoff

0.418

0.0169

0.347

N

on the ITF

0.091

0.266 

2.972

on the pickoff

0.561

0.044

0.484

Sorry for the delay of the start of the afteroon activity. 

Comments to this report:
nardecchia, pinto, gouaty, magazzu, tringali - 22:51 Friday 22 September 2023 (61755) Print this report

Today afternoon, it was not possible to reach LN2 most probably due to the unbalance of the B1p 56 MHz upper and lower sidebands (see fig.1).

In the lock started at 18.11 UTC (when CARM NULL 1F has been achieved), we added -11 as MICH offset in order to overimpose the sidebands  to proceed up to LN2, the trick worked (see fig.2) but the ITF unlocked at LN2 due to huge oscillation at ~55 Hz .  

Thus, to save the weekend, we decided to reverse the last step of DAS WP.

So, at 19.24 UTC, we came back to STEP_1S (61729) .

 

STEP_1S
19.24 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.090

1.857

on the pickoff

0.418

 0.0146

0.302

N

on the ITF

0.091

0.266 

2.972

on the pickoff

0.561

0.044

0.484

After that, the ITF achieved LN3 (at 20.25 UTC) without adding extra offsets (fig.3).

As requested by Romain, the BS TY SET was put to 0.

We left the ITF to Romain.

Images attached to this comment
Environmental Monitoring (Environmental Monitoring)
Paoletti, Tringali - 16:53 Friday 22 September 2023 (61752) Print this report
Comment to 48Hz WE F7 crossbar resonance high again (61751)
Images attached to this comment
Environmental Monitoring (Environmental Monitoring)
Paoletti, Fiori, Tringali, Pinto - 16:05 Friday 22 September 2023 (61751) Print this report
48Hz WE F7 crossbar resonance high again

As of mid-August, the 48.1 Hz that we label as "WE crossbar" became more excited. This occurred on Monday, August 14, within an unlock between 11:00 and 14:30 UTC. No particular action is reported in the logbook.

Today at 10:23 UTC we went to WEB and unplugged the mains cables from the motor crates, because we know that in the past there has been an unexpected coupling with some noise flowing through them.
The cables have been temporarily tagged (they are usually connected to a "smart power strip"), so that they can be reconnected in the right location.

We will test whether the 48 Hz noise is reduced at the next lock, after which we will make a final decision for the right system configuration.

It may be worth repeating the same action at NEB as well.

Images attached to this report
Comments to this report:
Paoletti, Tringali - 16:53 Friday 22 September 2023 (61752) Print this report
Images attached to this comment
Detector Operation (Operations Report)
menzione - 16:01 Friday 22 September 2023 (61750) Print this report
Operator Report - Morning shift

Upon arrival I found ITF LOCKING_IR and SUS+BS Metatron node stuck due to the problem on SUS_BS DSP board (BS gray flags on DMS). It prevented the lock, so I contacted Bersanetti who bypassed the BS top stage in the automation .
  - Experts allerted in case of BS_DSP Board reset or replacement: Passuello, Carbognani, Boschi, Ruggi, Bersanetti
At 08:45 UTC the lock acquisition restarted (COMMISSIONING MODE again set).

The planned TCS/ISC activity on differential DAS tuning (Nardecchia, Spinicelli) started at 09:00 UTC and went on without major problems except few unlocks.  

 

Sub-system reports

EnvMoni
12:16 UTC - LargeCoil_CEB process stuck -> stopped / restarted via VPM

Other
INF
two blackout occurred, several INF and ENV apparatus are still offline. Experts informed.

Downtime events

SUSP
(22-09-2023 06:00 - 22-09-2023 08:45) SUS_BS DSP board vs SUSP_BS Metatron node

Oncall events

SUSP
(22-09-2023 06:00 - 22-09-2023 08:00) Operator on site with expert from remote
Status: Ended
Description: SUS_BS DSP board investigation

Images attached to this report
AdV-COM (automation)
bersanetti - 10:33 Friday 22 September 2023 (61749) Print this report
BS top stage bypassed in the automation

I modified the code for the suspension nodes in order to skip the checks on the top stage; this applies to BS only. SUS_BS is operational again, there shouldn't be any other places where we try to access BS top stage.

Suspension Electronics (Suspension Electronic)
bersanetti - 9:38 Friday 22 September 2023 (61748) Print this report
Comment to SUS_BS DSP board investigation in progress (61747)

As deemed necessary, the North Arm is being kept locked by putting in pause the nodes SUS_BS, WARM_LOCK and ARMS_LOCK, and acting manually on NARM_LOCK.

Suspension Electronics (Suspension Electronic)
menzione - 9:07 Friday 22 September 2023 (61747) Print this report
SUS_BS DSP board investigation in progress

SUS_BS DSP board investigation in progress
ITF in TROUBLESHOOTING MODE

Comments to this report:
bersanetti - 9:38 Friday 22 September 2023 (61748) Print this report

As deemed necessary, the North Arm is being kept locked by putting in pause the nodes SUS_BS, WARM_LOCK and ARMS_LOCK, and acting manually on NARM_LOCK.

AdV-SGD (FDS commissioning)
zendri - 8:22 Friday 22 September 2023 (61746) Print this report
Comment to FIS injection with new B1 Photodiodes (61744)
At 5:35:55 UTC the squeezing injection was stopped (automation state "ready for FIS injection").
At 6:12 UTC the closing of the SQB1 shutter was completed.
Detector Operation (Operations Report)
menzione - 7:58 Friday 22 September 2023 (61745) Print this report
Operator Report - Initial status

ITF State: LOCKING_ARMS_IR
ITF Mode: LOCKING
Quick Summary: ISYS: in standard state, SBE loops closed, Susp loops closed, SUS_BS Metatron node stuck (TBC).
Activities ongoing since this morning: No activities ongoing

Images attached to this report
AdV-SGD (FDS commissioning)
vardaro, delaurentis, zendri - 1:52 Friday 22 September 2023 (61744) Print this report
FIS injection with new B1 Photodiodes

The aim of the shift was to check the performance of the FIS injection with the new B1 PDs

List of the activities:

As first action of the shift we measured the Squeezer parametric gain with the HD detector. We had 9mV max and 2.64 mV min  of HD 4 MHz mag. This correspond to a parametric gain of 3.41, i.e. 10.6dB o SQZ produced by the OPA

At 14:55 UTC we started to open the SQB1 Shutter and to rotate the SQB1 HWP2 half waveplate.

At 15:30 UTC with the automation in SQZ_INJECTED we closed the CC and AA loop and we realigned the SQZ beam toward the ITF

At 15:36 UTC we took the shot noise reference

At 15:44:30 UTC we started a phase scan. From the phase scan we measure roughly 6dB of ASQZ and less than 1dB of SQZ. See fig 1.

We found out that the angle of max SQZ level was 1.2 rad and the angle of max ASQZ level was 2.65 rad.

We measured ASQZ at 16:37 UTC and SQZ at 16:44 UTC. In figure 2 you can see in violet ASQZ, in BROWN shot noise taken at the end of the phae scan (16:10 UTC)  and in RED shot noise taken at 15:36 UTC. In green SQZ taken at 16:44 UTC. You can see that the level of the shot noise is different in both the case.

Thus we decided to take longer data stretches of Shot noise and SQZ:

  • figure 3: SQZ ingected up to  17 and 30, than shot noise. You can see that at the beginning the noise in hrec was higher and with a lot of fluctuations. Near the closure of the shutter it was better. When we closed the shutter you can see the step in the noise in Hrec. This is compatible with about 1 dB of SQZ. The cause of the noise was probably an high technical noise in the interferometer. It is also interesting that the 4MHz mag decreased by a factor 2%, but it is not enough to explain the high fluctuations at the beginning of the plot. Moreover are evident some correlation between the movement of SDB1 and the level of the  magnitude, and also some correlation between the coupling of the frequency noise in the ITF (SSFS) and the level of measured squeezing
  • figure 4: after the shot noise data we injected again squeezing. The level of the noise in hrec is similar to what we found at the end of the SQZ data in the first part of fig 3.
  • At 19 we injected ASQZ and we started to check the alignment, to be sure that the change in the B1 4MHz magnitude was not due to some clipping. We did the following acrions 
    • move SQB1 bench in horizontal X to see inf we could improve the level of the 4MHz mag on B1 (no effect)
    • move the Picomotors on M32 and M33 (small improvement)
    • move the retroreflector on SQB1 (small improvement)
    • optimize the SQB1 HWP1 to decrease the power of the spurious beam from the ITF on EQB1. This was done by looking the beat note of the LO and this beam on HD Detector (open LO Shutter, Put fast DAC offset at 20 mV, rotate SQB1 HWP2 to send all toward HD, improve SQB1 HWP1, come back with SQB1 HWP2)
    • At 20:21 UTC all these operation were concluded
  • Figure 5 SQZ (0.9dB at 18:45UTC), ASQZ (5.8 dB at 19:15 UTC), Shot (at 18 UTC).
  • Figure 6 the same of figure 5 in time domain

Some observations:

  • This time we injected 2dB of more SQZ wrt the last shift. We measure 2dB more of ASQZ but the same level of SQZ (maybe less). We had the impression that the technical noises at high frequency are higher than during the last shift
  • Figure 7. Comparison of B1 DC and B1 DC_NULL between new and old PDs. In the case of B1_DC null the noise is lower because the dark noise of PDs is lower, but B1_DC has the sane level. This seems due to the fact that the high frequency was not limited only by the dark noise of B1 PDs. Now the distance between B1_DC and B1_DC_NULL is higher wrt the older PDS, See fig 8.
  • We checked also the noise in Hrec during the last night, no SQZ inkected. And you can see that the level of hrec had oscillations compatible with an injection of 0.8 dB of SQZ (Fig 9). All these are probably evidence that we are still not shot noise limited.
  • When we removed the SQZ the sensitivity dropped of about 1 Mpc. Some effect is visible (fig 10)

We left the system in SQZ_INJECTED and the shutter open on SQB1. We are injecting SQZ all the night (fig11 preliminary data). At 21:15 UTC we closed the SLOW CC Loop (we forgot before) and we increase the CC gain from 20000 to 35000 (we did not update this value in SQZ_MAIN.ini file)

To remove SQZ injection put SQZ_MAIN in ready for FIS injection and close SQB1 shutter: (SQB1 pico, PM34_NC by plus (800000-1000000) steps)

Images attached to this report
Comments to this report:
zendri - 8:22 Friday 22 September 2023 (61746) Print this report
At 5:35:55 UTC the squeezing injection was stopped (automation state "ready for FIS injection").
At 6:12 UTC the closing of the SQB1 shutter was completed.
Detector Operation (Operations Report)
berni - 22:59 Thursday 21 September 2023 (61743) Print this report
Operator Report - Afternoon shift

The first part of the shift was dedicated to complete the DAS differential tuning.

At around 15:30 UTC started the planned activity on SGD; activity concluded around 20:20 UTC.

At 20:26 UTC I set the DQSTUDIES mode and then I engaged the autorelock.

Images attached to this report
AdV-COM (automation)
narnaud - 18:40 Thursday 21 September 2023 (61742) Print this report
Comment to Introduction of new DQSTUDIES ITF Mode & capability to trace actions from metamedm interfaces (61245)

Daily and weekly VIM summary plots now display DQSTUDIES segments.

AdV-TCS (CO2 laser projector)
mantovani, nardecchia, ruggi, magazzu - 17:19 Thursday 21 September 2023 (61741) Print this report
DAS differential tuning - PART 2

This morning, we proceeded with the DAS differential tuning. In order to try to evaluate the effect of each DAS ring, the NI outer ring has been decreased by 15 %. 

08.17 UTC: Locked at LN2

08.48 UTC: NI Outer ring power change (-15 % wrt the STEP_1S ) 

 

STEP2_S

 

08.48 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.078*(1+15/100)=0.090

1.615*(1+15/100)=1.857

on the pickoff

0.418

0.0127*(1+15/100)= 0.0146

0.263*(1+15/100)=0.302

N

on the ITF

0.091

0.219*(1-15/100)+0.08=

0.266 

1.172*(1-15/100)+1.8 =2.8

on the pickoff

0.561

0.036*(1-15/100)+0.013=

0.044

0.191*(1-15/100)+0.293=
0.455

09.13 UTC- 9.18 UTC: Manuel tested a new filter

09.18 UTC: ITF unlocked

09.25- 09.46 : Diego P. went to the CB on the platform.

At the next relock, the ITF working point appeared degraded (see fig.1).

Thus, in order to investigate the ratio between NI inner and outer ring, the NI Inner ring power has also been decreased by the same amount (-15 %) .

10.40 UTC:  NI Inner ring power change (-15 % wrt the STEP_2S ). 

 

STEP3_S

 

10.40 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.078*(1+15/100)=0.090

1.615*(1+15/100)=1.857

on the pickoff

0.418

0.0127*(1+15/100)= 0.0146

0.263*(1+15/100)=0.302

N

on the ITF

0.091

0.1861*(1-15/100)+0.08=0.238

1.172*(1-15/100)+1.8 =2.8

on the pickoff

0.561

0.026+0.013=0.039

0.191*(1-15/100)+0.293=
0.455

Also this step had a negative impact on the ITF WP, B1p was very high and we were not able to lock in low noise 2 due to unbalance of B1P PC 56 MHz.

Thus, with the ITF in CARM NULL 1F, the NI DAS inner and outer ring powers have been increased again, at 12.08 UTC.

 

STEP4_S

(= STEP1_S)

 

12.08 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.078*(1+15/100)=0.090

1.615*(1+15/100)=1.857

on the pickoff

0.418

0.0127*(1+15/100)= 0.0146

0.263*(1+15/100)=0.302

N

on the ITF

0.091

0.219*(1-15/100)+0.08=

0.266 

1.379*(1-15/100)+1.8 =2.972

on the pickoff

0.561

0.036*(1-15/100)+0.013=

0.044

0.225*(1-15/100)+0.293=

0.484

13.15 UTC: unlocked

13.24:WI DAS powers have been increased by 15%.

 

STEP5_S

 

12.39 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.090*(1+15/100)=0.104

1.615*(1+15/100)=2.13

on the pickoff

0.418

0.0146*(1+15/100)= 0.0169

0.302*(1+15/100)=0.347

N

on the ITF

0.091

0.219*(1-15/100)+0.08=

0.266 

1.379*(1-15/100)+1.8 =2.972

on the pickoff

0.561

0.036*(1-15/100)+0.013=

0.044

0.225*(1-15/100)+0.293=

0.484

At 14.05 UTC, the ITF locked in LN2.

It seems to be a good WP (see fig.2) but it is needed to dedicate some time to confirm it.   

We left these DAS powers identified as STEP 5_S for the afternoon shift. 

Images attached to this report
Detector Operation (Operations Report)
amagazzu - 15:58 Thursday 21 September 2023 (61740) Print this report
Operator Report - Morning shift

The shift was dedicated to the planned TCS/ISC, differential DAS tuning, carried out by Nardecchia and Mantovani. We found the ITF in LOW_NOISE_3 since yesterday night, at 6:00 UTC we attempted to go back to LOW_NOISE_2, but the ITF unlocked during the transition.
During the shift we were able to relock and reach LOW_NOISE_2 multiple times, from 11:05 UTC we set CARM_NULL_1F as target.
Activity still in progress.

Sub-system reports

DET
During a thermal transient and with the ITF in CARM_NULL_1F, the B1p Galvo loop opened due to saturation of the correction. Loop manually restored at 11:18 UTC and 11:40 UTC.

SBE
SQB2 Horizontal Loop opened at 5:34 UTC. At 6:00 UTC I closed it again via VPM, but the loop opened again at 8:10 UTC. From 8:57 UTC to 10:10 UTC I acted on the motors from 0 to 3 to restore the Horizontal and Vertical position of the bench close to its setpoints. Loop closed again at 10:09 UTC

SUSP
Since around 3:45 UTC of this night the loop signals related to BS ID, Vertical loop and related guardians became unavailable, while the loops remained closed. Thanks to the support of Carbognani we investigated the issue, and found that the board 41033 was not reachable anymore, preventing connection via DSP interface, SuspensionUI and Chassis Client.
Due to the impossibility of Boschi to intervene on the issue we contacted Passuello, who went to the Central Building after an unlock at 9:30 UTC to investigate.
Unfortunately the only action available is to restart the board, which would trigger an opening of BS ID: in case the connection is not restored, it would be impossible to close the loop again without a long maintenance intervention.
In agreement with the experts and the Commissioning Coordinator, we postponed the recovery of the connection.

Images attached to this report
Detector Characterisation (Broadband noise)
Tringali, Fiori - 15:09 Thursday 21 September 2023 (61728) Print this report
Hrec structures exicited during SIB1 bench shaking and near field magnetic injection

As for the SDB1 bench, we compared the Hrec structures excited during the SIB1 near-field sweep magnetic injections (#60083) and the noise injections on the SIB1 bench  (#61714).
Below, is a list of matched structures during the two noise injections.

Figure 1: Hrec spectrum during the magnetic (black curve) and mechanical noise injections (TY, TZ  red and yellow curves).

Figure 2, (20-70)Hz: 

  • a resonance structure is excited at ~28.1 Hz (TY dof for mechanical injection);
  • NE_MIR_permline_ext0 at 37.5 Hz and WE_MIR_permline_ext0 at 56.5 Hz  (TZ dof) are excited;
  • broad noise structure in the frequency region ~(60-65)Hz.

Figure 3, (70-140)Hz: 

  • a resonance structure is excited at ~76.8 Hz (TY dof);
  • a resonance structure is excited at ~101 Hz (TY dof) and 136.15 Hz.

Figure 4: SSFS_LF_line from ASC_pre | SSFS_LF_line from LSC_Acl_Moni @227.1 Hz and its sidebands ±2.09 Hz, ±4.29 Hz (slight excitation of TY dof) and +6.09 Hz are excited.
 

Images attached to this report
AdV-TCS (CO2 laser projector)
mwas - 15:02 Thursday 21 September 2023 (61739) Print this report
Comment to DAS differential tuning- CO2 (61729)

Figure 1 shows the OMC scan performed yesterday.

Figure 2 shows for comparison the last OMC scan I could find, done on Jul 24.

On figure 1 the carier order 2 mode is a bight higher (30mW vs 20mW in July). The order 3-6 modes look the same, modes 7 and 8 are now much brighter by about a factor 2.

Images attached to this comment
AdV-ISC (LSC Noise Budget)
mwas - 14:51 Thursday 21 September 2023 (61738) Print this report
Comment to RIN coupling in simulink noise budget (61731)

Given how well measurements match the model coupling at high frequency. I have replaced 6 month old measurement with the model in the noise budget, which will adjust the MICH offset in the simulation to match the RIN coupling measured at 1.5kHz.

Figure 1 shows the result of the projection during the ~20s when the PSTAB loop was open. Projection is good at high frequency, but under 100Hz there is still a factor few missing. To be sure what the coupling really is below 100Hz one would need to make a controlled measurement injecting RIN noise (without affecting too much the inteferometer control). The model correspond to a 1.2nm offset/RMS on MICH.

Figure 2 shows the projection during normal calm times, with a 0.7nm offset/RMS on MICH. The black curve correspond to PSTAB sensing noise, and the magenta curve to RF sideband RAM contamination of PSTAB sensing.

Figure 3 shows the corresponding input RIN to B1 RIN coupling used with a 0.7nm offset/RMS on MICH. The coupling comes out of an optickle simulation.

Images attached to this comment
Infrastructures (Air Conditioning)
cavalieri, soldani - 11:01 Thursday 21 September 2023 (61737) Print this report
NEB AHU PID Controller tuning

In order to reduce also NEB-AHU COLD loop oscillations, especialy associated with low temperature chiller regime, (according to WEB AHU PID Controller tuning), on September 13th, we modified the NEB-AHU PID Controller parameters, in detail:
14:35 UTC removed derivative term (from 90 sec to 0 sec) and increased proportional band from 3 to 4.
Fig.1 and Fig.2 summarizes consequences of our actions.
The removal of derivative term stops the oscillation (period 120sec). The residual oscillation in the correction signal, follows the chiller temperature (Fig2).
The reintroductiion of derivative term, 90 sec, on Sep 18th 8:30 UTC , in contrast with WEB, causes again loop instabilities (Fig3, Fig.4).

The derivative term has been removed on Sep 19th 13:58 UTC (Fig.5).


PID parameters:
Integration time: 360 sec
Derivative time 0 sec
Proportional band 4°

Images attached to this report
AdV-COM (AdV commissioning (1st part) )
masserot - 9:55 Thursday 21 September 2023 (61736) Print this report
Comment to PRCL_INPUT: B4_56MHz_{I,Q} weight trials (61723)

This morning the ITF was locked at LOW_NOISE_3 with 

  • PRCL_INPUT  =  72 *"B2_6MHz_I"  -120.062 * "B4_56MHz_I" + 38.373* "B4_56MHz_Q" .
  • The basic error signal for PRCL is B2_6MHz_I

With the ITF in LOW_NOISE_1 , for this lock period,  PRCL_INPUT was equal to

  • PRCL_INPUT  =  72 *"B2_6MHz_I"  -2.541 * "B4_56MHz_I" + 3.911* "B4_56MHz_Q"

At 2023-09-21-04h29m40-UTC  PRCL_INPUT was set be equal to only  72 *"B2_6MHz_I"  with a ramp of 120s:

  • the first plot show the trend of the full LOCK
  • the second plot is a zoom of the trend  when the  PRCL_INPUT weigths were updated
  • The third plot and its zoom  compare  the  ITF length error signals before(purple) and after (blue) the PRCL_INPUT channel weights update

The last plots show PRCL, MICH and SRCL  FFT Time  plots where the time period with PRCL_INPUT  equal to 62*B2_6MHz_I only is highlighted (red rectangle)

Images attached to this comment
Search Help
×

Warning

×